# Delayed takedown of illegal content on social media makes moderation ineffective

### v1

@@ Results
Based on these estimates, we run experiments with $10^{-4}\leq p\leq 10^{-2}$ (Values $p<10^{-4}$ would require a network with infeasibly large size, as discussed in Methods.) In our main analysis we report the reduction in illegal content prevalence, $1-\frac{\bar{I}_{\text{removal}}}{\langle\bar{I}_{\text{baseline}}\rangle}$. To show that this ratio is independent of $p$, we plot the scaling relationship between $\bar{I}_{\text{removal}}$ and $\bar{I}_{\text{baseline}}$ in Fig. 5a for different $p$ values from Table 2. We observe only a weak sublinear relationship, $\bar{I}_{\text{removal}}\sim\bar{I}_{\text{baseline}}^{0.94}$, suggesting that the ratio is nearly independent of $p$. We also plot in Fig. 5b the reduction in the prevalence of illegal content for different values of $p$ from Table 2. None of the differences between the values in our main results ($p=0.01$) and those for smaller $p$ values are statistically significant. These results suggest that our main results are robust with respect to the volume of illegal content in the system.
Figure 5: Robustness analysis for illegal content probability. (a) Scaling between illegal content prevalence in removal and baseline conditions. Each data point represents the proportion of illegal content in the system with and without removal. Results are based on 40 simulations for each value of $p$ from Table 2, using $\tau=2$. The gray line represents a least-squares fit of the log-transformed data, yielding a slope of $0.94\pm 0.01$. The vertical lines mark the illegal content probabilities from Table 2. (b) Illegal content reduction for different illegal content probabilities corresponding to the same countries and platforms, based on 10 simulations with $\tau=2$ (top) and $\tau=8$ (bottom). The result does not vary significantly ($P\geq 0.13$ across pairs of experiments with different $p$ values, using Mann–Whitney U tests with Bonferroni correction).
@@ Discussion
This paper examines empirical data on illegal content takedown on multiple social media platforms. Our agent-based model calibrated on this data reveals that removal of illegal content within a one-day window significantly reduces its prevalence and exposure, whereas more prolonged delays allow illegal content to permeate user feeds, diminishing the impact of the moderation. Delays beyond 46 days have no impact on the spread of illegal content.
@@
These findings have practical implications for social media moderation and regulation. By adopting more proactive takedown policies, platforms may reduce moderation delay and limit the spread of illegal content. Strengthening the use of and cooperation with trusted flaggers, who identify and report content violating community guidelines or legal standards, can further enhance content moderation. With human flaggers being a limited resource (29), efforts should be directed towards prioritizing the most time-sensitive cases.
@@
For legislation, the importance of prompt and timely moderation appears to support the argument for implementing tight illegal content removal deadlines, as imposed under NetzDG in Germany and KoPl-G in Austria prior to the DSA. However, the practical impact of such deadlines should be carefully assessed in regulatory debates for two reasons. First, regulatory takedown deadlines can only address the timespan between the notification of illegal content and its removal, not the period during which illegal content circulates in the network undetected. As a result, the impact of tight deadlines on the spread of illegal content is inherently limited. In fact, most platforms already report fast reactions after notifications: a median time below 24 hours according to 2024 DSA Transparency Reports and removal of 80% of illegal content within 24 hours according to 2022 national Transparency Reports for Germany and Austria.
@@
Second, tight takedown deadlines may reduce the time necessary for properly assessing the (un)lawfulness of content, increasing the risk of false moderation decisions and lowering the quality of content moderation. Flagged content requires careful verification, including report reviews and appeal resolutions, before removal decisions can be made. Any regulatory framework must ensure effective moderation of illegal content while remaining practical for platforms, and mindful of potential adverse effects. For example, in France, the Loi Avia was struck down by the Constitutional Council (30) due to concerns about restrictions on freedom of expression (31, 32). These considerations suggest that platforms should remove immediately only obviously illegal content. For cases requiring further assessment, platforms should take time for thorough review. Until the assessment is completed, platforms may implement alternative interventions to reduce the circulation of potentially illegal content.
Our model has several limitations. Certain aspects, such as the user activity levels, are calibrated on data from global Twitter users, while the DSA data focuses on European users. Unknown biases could stem from this combination of data sources. The model further presumes uniform spreading dynamics across various types of illegal content. In reality, the appeal of such content may vary widely; scams may attract a broader audience while human trafficking posts may be aimed at specific communities. Due to the lack of data on such content, our model does not distinguish among classes of illegal content or between illegal and legal content. If illegal content were more appealing, our findings would underestimate the adverse effects of time delay on moderation. For modeling content moderation, we make the simplifying assumption that when illegal content is removed, it is deleted from all user feeds. In practice, however, some content can be illegal only in certain jurisdictions and access can be selectively disabled for users in these countries. Our model does not consider such a differentiation.
The DSA-TDB has its own limitations. The database depends heavily on platform voluntary reporting, leading to inconsistencies across platforms. This reflects a fundamental weakness in self-reporting systems, as moderation decisions are often shaped by specific platform terms of service and community guidelines.
@@
Future research could extend our model by accounting for the fact that users often alter their posting patterns in response to moderation decisions (33, 34). The model could also compare the effects of takedown versus alternative types of interventions, as well as explore combinations of different intervention instruments. Finally, future models should consider the trade-offs between the harm caused by persistent illegal content and the unitended effects of moderation tools, e.g., the limitation of free speech caused by false-positive errors.
Methods
Model extensions
We extend SimSoM to model the heterogeneous activity of social media users, where most agents have low activity levels, and only a few are hyperactive. We model this by assigning an activity level $a$ to each user at the beginning of the simulation. We draw this activity from a power law $P(a)\sim a^{\gamma}$, where we obtain the exponent $\gamma=2.85$ by fitting empirical measurements of average daily posts (Fig. 1a). If $a<1$, the user (re)shares a single message with probability $a$ at each time step; else the user (re)shares $\lfloor a\rfloor$ messages.
We wish to model the probability of posting illegal messages by low- and high-risk accounts. Given the lack of empirical data about this probability, we used the posting of content from low-credibility sources as a proxy. Leveraging news source credibility ratings from NewsGuard (35), we categorized a sample of over 250 thousand Twitter users into a low-credibility and a high-credibility group by comparing the average ratings of sources they shared against two thresholds. We then inspected the distributions for posting low-credibility content by these two groups and found that both could be modeled using the parametric family of beta distributions. Based on this proxy, we model the probability $i$ of posting illegal messages by users in each group (low- and high-risk) with a beta distribution $P(i)\sim i^{\alpha-1}(1-i)^{\beta-1}$, where $\alpha>0$ and $\beta>0$ are parameters that express the mean $\frac{\alpha}{\alpha+\beta}$ and skewness of the distribution for that group.
Given the two group distributions, we can express the average illegal content probability across the full population as $p=s_{H}\frac{\alpha_{H}}{\alpha_{H}+\beta_{H}}+(1-s_{H})\frac{\alpha_{L}}{\alpha_{L}+\beta_{L}}$, where $s_{H}$ is the relative size of the high-risk groups, and the $\alpha$ and $\beta$ subscripts indicate the groups.
Since platforms typically do not provide access to original content, there is insufficient scrutiny to determine the legality of content and limited data on the share of illegal content (for an exception see Wagner et al. (36)). Hence, it is difficult to empirically calibrate the illegal content probability $p$ by estimating the percentage of illegal messages in the system. Therefore, we choose different values of $\alpha$ and $\beta$ to explore a broad range of values $10^{-4}\leq p\leq 10^{-2}$.
Content removal
The illegal content prevalence is defined as the average ratio of illegal content in circulation at time $t$: $I_{t}=\frac{1}{N}\sum_{i=1}^{N}f_{i,t}$ where $f_{i,t}$ is the fraction of illegal content in user $i$’s feed at time $t$ and $N$ is the number of agents.
For each set of parameters, we run 70 simulations without any content removal; these are averaged to obtain baseline metrics. We then run 70 simulations with content removal; the reported reduction is averaged across these runs. For example, the reduction in illegal content prevalence is calculated by $(\langle\bar{I}_{\text{baseline}}\rangle-\bar{I}_{\text{removal}})/\langle\bar{I}_{\text{baseline}}\rangle$, where $\bar{I}$ is the illegal content prevalence at the end of a simulation and $\langle\bar{I}_{\text{baseline}}\rangle$ is the baseline average. All simulations start from random conditions.
Data sources
More detailed data, suited for our purpose, is available in the newly introduced DSA Transparency Database, which compiles Statements of Reasons (SoRs) explaining single moderation decisions to affected users. These SoRs include the reasons for moderation, as well as information on dates of posting and moderation decisions. We therefore use this data source to estimate the empirical half-life ($\tau$) values. We collect SoRs from the DSA-TDB between January 01 and December 31, 2024. We capture relevant records using the strictest filter, where the platforms explicitly state that the moderation is due to illegal content. [^1] We select SoRs that have DECISION_GROUND_ILLEGAL_CONTENT in the decision_ground field. The database also includes the value DECISION_GROUND_INCOMPATIBLE_CONTENT for the same field. This can cover content that is illegal, which would be indicated in an optional boolean field incompatible_content_illegal. Since platforms do not use this reporting category consistently, we did not include it in our dataset for calibration. We consider four major social media platforms with a sufficient number of relevant SoRs: Facebook, Instagram, YouTube, TikTok, and Snapchat. We exclude LinkedIn and Pinterest due to a lack of SoRs related to illegal content and Twitter/X due to suspicious patterns revealed in its DSA data, as also highlighted in prior work (37). This results in 8,613 SoRs submitted by Facebook, 3,005 by Instagram, 408,400 by YouTube, 93,358 for TikTok, and 2,142 by Snapchat. The time delay in illegal content removal is defined as the difference between the moderation date (application_date field) and the content posting date (content_date field).
Follower network
Simulations are conducted on a follower network derived from empirical Twitter data. The network connects users obtained from a 10% random sample of public tweets (38). The dataset comprises users who posted links to news articles, with likely automated accounts excluded. The original follower network constructed from this dataset has $58{,}296$ nodes and $10{,}632{,}023$ edges. We reduce the size of this network to speed up the simulations in our main experiment (Section Effects of time delay). We employ $k$-core decomposition to identify $N=10{,}006$ nodes that constitute the $k=94$ core. Subsequently, a random selection of edges is removed to match the density of the original network. This procedure produces $E=1{,}809{,}798$ edges, with each node having an average of roughly 180 friends or followers.
For the robustness analyses varying high-risk group size and illegal content probability (Section Robustness), we need a large follower network to run meaningful scenarios using low values of these parameters ($s_{H}=10^{-3}$ and $p=10^{-4}$). Therefore we use the original network, but remove a random set of edges to speed up the simulations. The average in/out-degree is thus reduced to 18 friends per node, resulting in a network with $N=58{,}296$ nodes and $E=1{,}063{,}202$ edges.
To test the robustness of our results with respect to different network structures, we create a third, synthetic network with features ubiquitous in real-world networks, namely, the presence of hubs and clustering (directed triads). To this end, we use a directed variant of the random-walk growth model (39). We configure the parameters of this model to generate a network comparable in size and average degree to the network used in the main experiment. In particular, we initialize the network with 181 fully connected nodes and assume new nodes have fixed out-degree $k_{out}=180$. New nodes are added by first randomly selecting one node to follow. For the remaining $k_{out}-1$ connections, each new node has an equal probability to follow a random friend of the initially chosen node or a randomly selected node in the network. This results in a network with $N=10{,}006$ nodes and $E=1{,}810{,}905$ edges; each node has on average approximately 180 friends/followers.

### v2

@@ Results
Second, we wish to explore how the main result is affected by the volume of illegal content, which varies across platforms and countries. To estimate a plausible range for $p$, let us analyze the Future of Free Speech policy report (34). The report includes comment data collected from Facebook and YouTube across Germany, France, and Sweden during two weeks in June–July 2023. It also includes the fraction of comments that disappear ($\frac{\text{disappear}}{\text{total}}$) and the proportion of disappeared comments that are legal ($\frac{\text{legal}}{\text{disappear}}$). From this, we estimate the fraction of comments that are illegal: $\frac{\text{illegal}}{\text{total}}=\frac{\text{disappear}}{\text{total}}\times\frac{\text{illegal}}{\text{disappear}}=\frac{\text{disappear}}{\text{total}}\times(1-\frac{\text{legal}}{\text{disappear}})$. Table 1 shows that this apprach yields a range between 0.00002–0.009. We run experiments using these estimates (except for the smallest value, which would require a network with infeasibly large size, as discussed in Methods). In our main analysis we report the reduction in illegal content prevalence, $1-\frac{\bar{I}_{\text{removal}}}{\langle\bar{I}_{\text{baseline}}\rangle}$. To show that this ratio is independent of $p$, we plot the scaling relationship between $\bar{I}_{\text{removal}}$ and $\bar{I}_{\text{baseline}}$ in Fig. 5a for different $p$ values from Table 1. We observe only a weak sublinear relationship, $\bar{I}_{\text{removal}}\sim\bar{I}_{\text{baseline}}^{0.94}$, suggesting that the ratio is nearly independent of $p$. We also plot in Fig. 5b the reduction in the prevalence of illegal content for different values of $p$ from Table 1. We observe more noise for small $p$, but none of the differences between the values in our main results ($p=0.01$) and those for smaller $p$ values are statistically significant. These analyses suggest that our main results are robust with respect to the volume of illegal content in the system.
Figure 5: Robustness analysis for illegal content probability. (a) Scaling between illegal content prevalence in removal and baseline conditions. Each data point represents the proportion of illegal content in the system with and without removal. Results are based on 40 simulations using $\tau=2$. Each simulation is done with an empirical estimate of $p$ (vertical lines) for Facebook and YouTube users from Germany, France, and Sweden, derived from the data reported in Table 1. Estimates of $p$ are based on the proportion of disappeared content that is subsequently judged to be legal. The gray line represents a least-squares fit of the log-transformed data, yielding a slope of $0.94\pm 0.01$. (b) Illegal content reduction for different illegal content probabilities corresponding to the same countries and platforms, based on 10 simulations with $\tau=2$ (top) and $\tau=8$ (bottom). The result does not vary significantly ($P\geq 0.13$ across pairs of experiments with different $p$ values, using Mann–Whitney U tests with Bonferroni correction).
@@ Discussion
This paper reveals how delays in illegal content takedown impact its effectiveness. Using empirical takedown data for five social media platforms from the DSA-TDB, we find large variations in average takedown delays. While platforms like TikTok typically respond within a week, illegal content often remains online for over a month on platforms like Facebook, Instagram, Snapchat, and YouTube. Our agent-based model calibrated on this data shows that long delays can substantially diminish the impact of content removal. Although this relationship may appear intuitive, our analysis reveals non-linear dynamics between takedown speed and moderation effectiveness, including critical delay thresholds. Specifically, we show that removing illegal content within one day significantly reduces its prevalence and exposure, whereas prolonged delays allow illegal content to permeate user feeds — beyond 46 days, takedown becomes futile in curbing the spread of illegal content.
@@
These findings have practical implications for social media moderation and regulation. By adopting more proactive takedown policies, platforms may reduce moderation delay and limit the spread of illegal content. Another approach to improving content moderation is to strengthen collaboration with trusted flaggers — individuals or organizations that identify and report content violating community guidelines or legal standards. With human flaggers being a limited resource (35), efforts should be directed towards prioritizing the most time-sensitive cases.
@@
For legislation, the importance of timely moderation appears to support the argument for implementing tight illegal content removal deadlines, as imposed under NetzDG in Germany and KoPl-G in Austria prior to the DSA. However, the practical impact of such deadlines should be carefully assessed in regulatory debates for two reasons. First, regulatory takedown deadlines can only address the timespan between the notification of illegal content and its removal, not the period during which illegal content circulates in the network undetected. Therefore, the impact of tight deadlines on the spread of illegal content is inherently limited. In fact, most platforms already report fast reactions after notifications: a median time below 24 hours according to 2024 DSA Transparency Reports and removal of 80% of illegal content within 24 hours according to 2022 national Transparency Reports for Germany and Austria. These figures suggest that a large portion of the long delays measured from the DSA-TDB data occurs before illegal content is detected or flagged. To better inform policy, platforms should be required to report not only the timestamps for publication and removal of content, but also the detection/notification timestamps to the DSA Transparency Database. This would allow to identify how long content has circulated in the network undetected as well as the moderation delay in response to notifications.
@@
Second, tight takedown deadlines may reduce the time necessary for properly assessing the (un)lawfulness of content, increasing the risk of wrong moderation decisions and thus lowering the quality of content moderation. Flagged content requires careful verification, including report reviews and appeal resolutions, before removal decisions can be made. Any regulatory framework must ensure effective moderation of illegal content while remaining practical for platforms, and mindful of potential adverse effects. For example, in France, the Loi Avia was struck down by the Constitutional Council (36) due to concerns about restrictions on freedom of expression (22, 23). These considerations suggest that platforms should remove immediately only obviously illegal content. For cases requiring further assessment, platforms should take time for thorough review. Until the assessment is completed, platforms may implement alternative interventions to reduce the circulation of potentially illegal content.
While our findings are robust, several limitations warrant consideration. The lack of empirical data on illegal content results in several assumptions about the content production process in our model. First, the model focuses on the spread of content already posted, without accounting for the user incentives or legal risks influencing decisions to post illegal material. We assume static probability distributions of sharing such content, though in reality, this behavior may evolve in response to moderation policies, potentially declining under stricter or faster moderation. Second, as mentioned in Methods, we assume uniform appeal and spreading potential across all types of content. If illegal content is generally more engaging, the model may underestimate the impact of delayed takedown, and vice versa. Lastly, we triangulate data from multiple sources, which may introduce inconsistencies. For instance, user activity is calibrated using global Twitter data, whereas the DSA database focuses on European users. This mismatch could lead to unknown biases in our calibration.
Our model of the moderation process also has limitations. The DSA-TDB data only captures the downstream outcomes of content moderation, neglecting the detection process. Our model thus treats all illegal content as equally detectable, which may overlook important variation in the detection of different content types. We also make the simplifying assumption that when illegal content is removed, it is deleted from all user feeds. In practice, however, some content may be illegal only in certain jurisdictions and access may be selectively disabled for users in these countries. Our model does not consider such a differentiation.
The database itself has limitations. It relies heavily on voluntary reporting by platforms, leading to inconsistent coverage and reporting quality both across and within platforms (31). This highlights a broader issue with self-reporting systems. Moderation decisions are shaped by platform-specific policies and guidelines, which may not always align with legal standards: content labeled as “illegal” may in fact be lawful but removed due to internal platform rules. The issue of false positives may have inflated the inferred probability of illegal content from the DSA-TDB.
@@
Future research could extend our model by accounting for the fact that users often alter their posting patterns in response to moderation decisions (37, 38). One could also extend the model to compare the effects of takedown versus alternative types of interventions, as well as explore combinations of different intervention instruments. Finally, future models should consider the trade-offs between the harm caused by persistent illegal content and the unitended effects of moderation tools, e.g., the limitation of free speech caused by false-positive errors.